Ticket #599 (closed task: fixed)

Opened 13 years ago

Last modified 11 years ago

[M] Set up regular harvesting from all DPs (currently on glue)

Reported by: selatham Owned by: sdonegan
Priority: required Milestone: Reporting
Component: discovery Version:
Keywords: Cc:

Description (last modified by selatham) (diff)

  • NOCS works
  • PML/NEODAAS works
  • BODC - works
  • BADC - works
  • NEODC - works
  • BGS - Rachel Heaven. Works. They have an OAI on a public machine. Done harvested and ingested 227 records.
  • CEH - Nicolas Bertrand. they have set up an oai with 1 record. Harvested successfully. This ingests OK - just about(date problems). What about their numerous records in the old system?
  • BAS have records in an OAI but are invalid so can't ingest.
  • NCAR have nothing in their OAI at the moment. Ingested their old records. Some fail on parsing, but at least 42 visible.
  • WDCC records have &nbsp problem when harvesting. Ingested their old DIF records OK. These appear OK in discovery, links work.
  • TPAC six records harvested and ingested via OAI.
  • UKHO(MDIP) having &nbsp problem, so not turned on.
  • DASSH(MDIP) harvests OK, ingested, but not regular.

Change History

comment:1 Changed 12 years ago by selatham

  • Status changed from new to assigned
  • Description modified (diff)
  • Milestone changed from PreBeta to MeetingsJan2007

NEODC done.

comment:2 Changed 12 years ago by lawrence

  • Milestone changed from MeetingsJan2007 to System Integration

comment:3 Changed 12 years ago by selatham

  • Milestone changed from System Integration to PROD

Code etc is there just turned OFF on Dev system currently. Production system not up yet.

comment:4 Changed 12 years ago by selatham

  • Milestone changed from PROD to PROD Final

comment:5 Changed 12 years ago by selatham

  • Milestone changed from PROD Final to Replace Metadata Gateway

DEV - Nightly harvesting is now turned on for NOCS, BODC, BADC, NEODC and the MDIP partners DASSH, UKHO.

I'm holding off on PML(neodaas) until Mike tells me that their metadata production pipe is fixed.

LIVE needs NCAR, WDCC, BGS, BAS, CEH also.

comment:6 Changed 12 years ago by selatham

  • Description modified (diff)
  • Summary changed from [M] Set up regular harvesting from all DPs. to [M] Set up regular harvesting from all DPs (currently on glue)
  • BGS - Rachel Heaven. Works. They have an OAI on a public machine. Done harvested and ingested 227 records.
  • CEH - Nicolas Bertrand. they have set up an oai with 1 record. Harvested successfully. This ingests OK - just about(date problems).
  • BAS have records in an OAI but are invalid so can't ingest.
  • NCAR have nothing in their OAI at the moment. Trying to ingest their old records.
  • WDCC records have &nbsp problem when harvesting. Trying to ingest their old DIF records.
  • TPAC six records ingested, but OAI disappeared.

comment:7 Changed 12 years ago by selatham

  • Description modified (diff)

comment:8 Changed 12 years ago by selatham

  • Description modified (diff)
  • NCAR have nothing in their OAI at the moment. Ingested their old records. Some fail on parsing, but at least 42 visible.

comment:9 Changed 12 years ago by selatham

  • Description modified (diff)

comment:10 Changed 11 years ago by lawrence

  • Status changed from assigned to new
  • Owner changed from selatham to sdonegan

can we split this into tickets per harvest target? then you can close this as a duplicate

comment:11 Changed 11 years ago by lawrence

Note that ticket:840 includes UKHO, which I've closed as a duplicate even though you'll probably have to open up a new one ...

comment:12 Changed 11 years ago by lawrence

See also ticket:317!

comment:13 Changed 11 years ago by sdonegan

  • Status changed from new to closed
  • Resolution set to fixed
Note: See TracTickets for help on using tickets.